OpenStudy (aroub):

Please integrate cos^2x showing steps :)

OpenStudy (owlfred):

\[\cos^2 x = \frac{1 + \cos(2x)}{2} \]

OpenStudy (aroub):

Yess! I didn't get this part. Cos^2x is equal to that^??

OpenStudy (owlfred):

Yes it is. Here is how you derive it.\[\cos^2 x - \sin^2 x = \cos(2x)\]\[\Rightarrow \cos^2 x - (1 - \cos^2 x) = \cos(2x)\]\[\Rightarrow 2\cos^2 x - 1=\cos(2x)\]\[\Rightarrow \cos^2 x = \dfrac{1 + \cos(2x)}{2}\]

OpenStudy (aroub):

Ohhhh

OpenStudy (aroub):

THANK YOUUU :)

OpenStudy (owlfred):

No problem. Welcome back to OS!

OpenStudy (aroub):

Haha, thanks!